Check whether the function, f , defined below, is uniformly continuous or not:
f (x) =x^1/2 , x belong to [1,2].
1
Expert's answer
2020-04-19T16:39:46-0400

Consider the function "f(x)=\sqrt{x}"


The domain of the function "f(x)=\sqrt{x}" is "[0,\infty)"


Here, the given interval is "[1,2]" which lies within the domain of the function "f(x)=\sqrt{x}"


So, the function is uniformly continuous in the interval "[1,2]" including end points as well.


Therefore, the function "f(x)=\sqrt{x}" is uniformly continuous in the interval "[1,2]".




As shown in the graph, the function is continuous in the interval "[0,\infty)" , so function will be uniformly continuous in any interval lies within this domain.